OpenStudy (anonymous):

for the function f(x)=(3-4x)^2, find f^-1. Determine whether f^-1 is a function.

OpenStudy (anonymous):

To find the inverse, where there is x, put y. Set new expression equal to x, then solve for y.

OpenStudy (anonymous):

is this correct? sq rt y^-3/-4=f^-1(x) and it is not a function

OpenStudy (anonymous):

No. Try my method above, first punch in y where there is x.

OpenStudy (anonymous):

what is y is it -1?

OpenStudy (anonymous):

((3-4(-1))^2

OpenStudy (anonymous):

(3+4)^2=49

OpenStudy (anonymous):

Not -1, the actual letter y, so you have x=(3-4y)^2. Now solve for y

OpenStudy (anonymous):

I have 16y^2-24y+9=x

OpenStudy (anonymous):

You are going the wrong route. From my equation above, find the square root of both sides. That way it is cleaner, you have only one y value. Solve for y.

OpenStudy (anonymous):

sq rt x^-3/-4=y???

OpenStudy (anonymous):

I don't know where you are getting these numbers from. Take it one step at a time.

OpenStudy (anonymous):

I am squaring each side and I have sq rt x=3-4y . Is this the first step?

OpenStudy (anonymous):

Yes now, just solve for y. Easy.

OpenStudy (anonymous):

easy for you. step 2: sq rt x-3=-4y: step 3 (sq rt x-3)/-4=y.

OpenStudy (anonymous):

in solving this, first move the minus 4y on the other side to move the negative. Your answer is correct if you move the - from in front of the 4.

OpenStudy (anonymous):

is this a function?

OpenStudy (anonymous):

You have to punch it in your TI-84 graph and look at the table

OpenStudy (anonymous):

great-I don't know how to do that on the calculator

OpenStudy (anonymous):

To be a function, each x-value must go to only one y value. Looks like that happens.

OpenStudy (anonymous):

thanks for your help and explanations